OpenStudy (anonymous):

evaluate f(3) for f(x)=-x^2+2x, x<1 -2x+3, x>1

OpenStudy (amigatour):

(3)=-x^2+2x, x<1 f(3)=−(3)^2+2(3) f(3)=-9+6= -3 -3<1 The other one is false because -3 is not greater than 1. Here is why: f(3)= -2x+3, x>1 -2(3)+3= -3 -3>1 is not true.

OpenStudy (anonymous):

OK. That's what I sort of thought but wasn't sure.